### v2

# Zoom-IQA: Image Quality Assessment with Reliable Region-Aware Reasoning

Abstract
Image Quality Assessment (IQA) is a long-standing problem in computer vision. Previous methods typically focus on predicting numerical scores without explanation or providing low-level descriptions lacking precise scores. Recent reasoning-based vision language models (VLMs) have shown strong potential for IQA by jointly generating quality descriptions and scores. However, existing VLM-based IQA methods often suffer from unreliable reasoning due to their limited capability of integrating visual and textual cues. In this work, we introduce Zoom-IQA, a VLM-based IQA model to explicitly emulate key cognitive behaviors: uncertainty awareness, region reasoning, and iterative refinement. Specifically, we present a two-stage training pipeline: 1) supervised fine-tuning (SFT) on our Grounded-Rationale-IQA (GR-IQA) dataset to teach the model to ground its assessments in key regions, and 2) reinforcement learning (RL) for dynamic policy exploration, stabilized by our KL-Coverage regularizer to prevent reasoning and scoring diversity collapse, with a Progressive Re-sampling Strategy for mitigating annotation bias. Extensive experiments show that Zoom-IQA achieves improved robustness, explainability, and generalization. The application to downstream tasks, such as image restoration, further demonstrates the effectiveness of Zoom-IQA.
Figure 1: (Upper) Current IQA methods are non-interactive, leading to inferior assessments. They either spot only partial flaws (e.g., slightly overexposed or slightly blurred) or make factually incorrect claims (clear and well-lit), resulting in erroneous judgments. Our Zoom-IQA uses interactive, region-aware reasoning: it first hypothesizes flaws (green text), then grounds them by cropping (orange text), and finally verifies the degradation (blue text). This hypothesize-and-verify loop provides a complete and accurate assessment. (Lower) Our model’s reasoning outputs also benefit downstream tasks, such as text-guided image restoration with SUPIR [75]. Our prompt enables a far superior restoration compared to those guided by other IQA methods or SUPIR’s default VLM, LLaVA-1.5-13b [32].
11footnotetext: Corresponding author
1 Introduction
Image Quality Assessment (IQA) is a fundamental task in computer vision, aiming to evaluate the perceptual quality of images in alignment with human perception. Its importance has grown rapidly as IQA models increasingly serve as critical perceptual reward signals for improving modern algorithms. Specifically, IQA serves as a component in frameworks like Reinforcement Learning (RL) from Human Feedback (RLHF) [46, 14, 63, 17] to align outputs with human preferences. Similarly, in image restoration, IQA scores are used as differentiable rewards [55] or in Direct Preference Optimization (DPO) [66, 4] to guide models toward perceptually superior results.
The emergence of vision language models (VLMs) like CLIP [44] opened a promising direction for IQA [59], further advanced by large-scale VLMs [58, 32, 2] that leverage broad knowledge to better align with human perception. Existing VLM-based IQA methods fall into two categories: (1) Score-based methods (e.g., Q-Align [65] and DeQA-Score [71]), which emphasize accurate scores but lack the ability to provide textual descriptions; and (2) Description-based methods (e.g., DepictQA series [73, 72]), which offer detailed explanations but rely on SFT data with descriptions generated from ground-truth synthetic distortions. To bridge this gap, recent methods like Q-Insight [29] and VisualQualityR1 [67] introduce RL to unify quality scoring and textual reasoning using only score labels as rewards.
Despite impressive capabilities, these VLM-based IQA methods remain non-interactive. They generate responses in a single pass without any mechanism for iterative visual refinement or correction. As highlighted in complex visual tasks (e.g., object detection [36, 50] and visual question answering (VQA) [74, 38]), the lack of intermediate visual grounding may lead to unreliable responses and reasoning, especially under complex scenarios. Similarly, the nature of IQA requires subjects to interpret complex images by “zooming in” on key regions. The importance of this behavior is highlighted by the DiffIQA dataset [8], which provided annotators with a zoom-in feature to inspect details. Such a region-aware interaction plays a helpful role in understanding the quality of an image, but is absent in existing VLM-based IQA models, confining their reasoning to the text domain and limiting effective use of visual information, as shown in Fig. 1.
Enabling a VLM-based IQA method to dynamically “crop and zoom” for iterative region-aware assessment faces two core challenges. (1) Region-aware Learning. The model must learn where to focus and how to transform regions (e.g., crop, zoom) based on its own partial textual deliberations, similar to the grounding in VQA [79, 48]. However, grounding IQA is more challenging than VQA. In VQA, grounding is often explicit, i.e., an answer can be tied to discrete, localizable semantic objects, for which extensive annotations are available [79, 48]. In contrast, an IQA score is a holistic judgment aggregated from numerous, complex factors. The core difficulty is the lack of supervision specifying which regions a human prioritized to arrive at their final score. While recent IQA grounding datasets [6, 9] provide static distortion masks, these methods fail to reveal the criticality of those regions to the overall human assessment or capture the dynamic reasoning path that led to the final judgment. (2) Self-Guided Reasoning Policy. The model must learn a dynamic policy on when to trigger detailed visual inspection rather than relying on random exploration or preset rules for zooming. This requires an unsupervised iterative cognitive process, i.e., performing a holistic assessment, identifying its own uncertainty about a specific region, and only then deciding to “zoom in” for refinement.
To bridge such gaps, we make two primary contributions. First, we introduce Grounded-Rationale-IQA (GR-IQA), a fine-grained dataset curated to facilitate the development of interleaved text-image Chain-of-Thought (CoT) reasoning. Directly harnessing advanced VLMs such as Gemini [12] to label IQA data with reasoning and scores can suffer from misalignment between visual inputs and reasoning outputs due to hallucination [27, 34]. Our GR-IQA is designed to avoid such hallucination by providing rationales that are verifiably grounded in visual regions. Specifically, our curation pipeline introduces two key modules: (1) Visual Reliance Filtering (VRF), which enforces grounding by measuring the generative output shift (with vs. without the image); and (2) Hint-Augmented Consistency Filtering (HACF), which filters hallucination-like descriptions at the sentence level to maintain fluency. Second, we propose Zoom-IQA (Zoomable Region Reasoning for Reliable Image Quality Assessment), a novel framework designed to enhance the reasoning reliability of IQA with region awareness. Zoom-IQA is trained in two stages: it first learns formatted grounding (how to “zoom”) via supervised fine-tuning (SFT) on our GR-IQA dataset, and then learns a dynamic policy (when to “zoom”) via reinforcement learning (RL). Such learned policy allows Zoom-IQA to operate iteratively, moving beyond “single-pass” methods to identify uncertainty and refine its assessment, achieving truly interactive visual reasoning. To stabilize the training process, we further propose the KL-Coverage regularizer, designed to prevent a collapse in reasoning path diversity, which often leads to a severe “mode collapse” in predicted scores. A Progressive Re-sampling Strategy is also developed to mitigate bias from imbalanced annotations.
Our Zoom-IQA is evaluated across diverse datasets and IQA tasks, demonstrating superior performance over both conventional IQA metrics and recent SFT-driven large language models. Moreover, Zoom-IQA exhibits impressive zero-shot generalization, such as effectively guiding image restoration models at test time, which highlights the robustness and real-world applicability of its region reasoning.
Figure 2: An overview of our two-stage framework. Stage (1), Grounded Quality Rationale Learning (Sec. 3.1), first uses SFT to teach the model how to correctly execute the crop action. Stage (2), Self-Guided Exploration (Sec. 3.2), then uses RL to let the model learn what to crop, allowing it to discover regions that lead to a deeper understanding of image quality.
2 Related Work
Image Quality Assessment. Previous IQA works are broadly divided into full-reference (FR) and no-reference (NR) approaches, based on the availability of a pristine reference image. As our work does not require a reference, we focus on the more challenging NR-IQA task. Conventional NR-IQA methods [37, 39, 40, 41, 42] relied on hand-crafted, degradation-aware features to predict the final quality score. Subsequent deep learning-based models [24, 3, 53, 25, 11, 54, 35, 43, 80] replaced this pipeline, directly predicting quality scores using end-to-end trainable neural networks. Nonetheless, these models often suffer from significant performance degradation on out-of-distribution (OOD) data, limiting their practical usage and generalizability.
Vision Language Models in Image Quality Assessment. Vision language models (VLMs) [44, 58, 32, 2] have been extensively studied in image quality assessment (IQA), leveraging their powerful cross-modal understanding and strong generalization capabilities. These works typically focus on one of two objectives: providing numerical quality scores [59, 65, 71, 81, 33] or generating visual quality descriptions [64, 73, 72, 9]. Specifically, CLIP-IQA proposes to harness CLIP [44] for image quality assessment from multiple aspects. DOG-IQA [33] attempts to mimic the human evaluation process (e.g., zooming in to evaluate specific areas). However, these models lack the necessary reasoning capabilities and cannot dynamically decide which regions to inspect. A common way to alleviate such limitations is via pre-processing, i.e., using a pre-trained segmentation model to crop sub-images and then computing the final score as a weighted average of these crops. Recent works like Q-Insight [29], VisualQuality-R1 [67], and Q-Ponder [5] propose to employ reinforcement learning (RL) to leverage the reasoning capabilities of VLMs, enabling image quality rating as well as textual justifications. However, the reasoning chains generated by these methods remain purely textual. They lack dynamic interaction with the image (e.g., cropping or zooming) to evaluate specific regions—a process crucial to human assessment. Consequently, visual evidence is insufficiently explored, and the models fail to ground their textual reasoning in verifiable image regions, limiting both the reliability and interpretability of their outputs.
Vision Language Models with Multimodal Reasoning. Recent advances in enhancing the reasoning capabilities of VLMs have significantly improved the performance of VLMs on challenging tasks, such as mathematical problem solving [38, 19, 76], VQA [74, 38], and object detection [36, 50, 74]. However, these models typically generate reasoning chains composed solely of natural language. This text-only reasoning can be opaque and often lacks sufficient grounding in the visual input’s fine-grained details. To address this, recent works in VQA [79, 78, 51, 21] propose to integrate evidence regions into the reasoning process. By equipping VLMs with capabilities like iterative zoom-in and region-of-interest selection, these methods demonstrated boosted performance. Improved interpretability is further gained via visual-linguistic interactive reasoning. However, these VQA methods generally rely on heavily annotated evidence regions (e.g., bounding boxes) for training. Such fine-grained regional labels and corresponding reasoning trajectories are critically expensive and lacking in the IQA domain.
3 Methodology
Our method, Zoom-IQA, is trained via a two-stage pipeline (illustrated in Fig. 2): 1) Supervised Fine-Tuning for Grounded Quality Rationale Learning: We first leverage our GR-IQA dataset to teach the VLM the foundational “how-to” skills: grounding textual rationales in visual regions and executing the “zoom” action (Sec. 3.1). 2) Reinforcement Learning for Self-Guided Exploration: To enable iterative refinement without exhaustive supervision, we employ Reinforcement Learning to derive a dynamic policy that optimizes the deployment of these skills. Specifically, we adopt Group Relative Policy Optimization (GRPO) [49] as our optimization framework, facilitating stable and efficient policy exploration (Sec. 3.2).
3.1 Grounded Quality Rationale Learning
Recalling the challenges from the Introduction, grounding IQA is uniquely challenging. While VQA supervision can link answers to semantic regions, IQA supervision is often limited to static distortion masks. This static approach fails to capture the dynamic reasoning path of how or why specific regions influence the final holistic assessment.
3.1.1 Grounded-Rationale-IQA (GR-IQA) Dataset.
To bridge this gap, we curate the Grounded-Rationale-IQA (GR-IQA) dataset with approximately 7,000 reasoning trajectories. This curation is performed through a novel pipeline (Fig. 3) that features our two key modules: 1) Visual Reliance Filtering (VRF); 2) Hint-Augmented Consistency Filtering (HACF).
Data Generation. We prompt the closed-source VLM, Gemini-2.5-pro [12], on KonIQ dataset [18] images using a structured prompt. This comple the VLM to generate a two-part response: a textual rationale (within <think>) and a JSON action (within <answer>). The textual rationale is strictly constrained to a four-part format: (1) a holistic Image Quality Summary; (2) Directions for Improvement; (3) Issues to Avoid; and (4) a Decision & Rationale where the model must decide if the image is an “easy case” (requiring a “final” tool) or a “hard case” (requiring a “crop” tool) and justify its choice. The <answer> block then contains the chosen “tool” (“final” or “crop”), the rating, and a conditional “bbox”. This structured process encourages the VLM to link its score to regional evidence and to self-assess uncertainty (i.e., “zoom” or not). This raw output forms the data for our GR-IQA dataset, which is then passed to our filtering modules (VRF and HACF).
Visual Reliance Filtering (VRF). Since VLMs inherit strong language modeling capabilities from LLMs, they may over-rely on textual co-occurrence patterns, leading to hallucinated outputs [34, 27]. To address this, we propose Visual Reliance Filtering (VRF). This module filters low-reliance samples by comparing the VLM’s output in two distinct scenarios: 1) conditioned on both the image $I$ and the textual rationale $R_{A}$, and 2) conditioned only on the rationale $R_{A}$. If the outputs are too similar, we discard the sample, as this indicates the visual input $I$ was not essential and the answer could be reproduced from the text alone.
Hint-Augmented Consistency Filtering (HACF). While VRF ensures the final <answer> is grounded, the textual rationale $R$ (the <think> block) may still contain unfaithful statements. To filter these low-veracity samples, we employ a powerful LLM, Qwen-2.5-32b [69], denoted $\text{LLM}_{\text{Rater}}$, to perform a holistic assessment of the entire rationale $R$. Crucially, to aid this judgment, the $\text{LLM}_{\text{Rater}}$ is provided not only with the full rationale $R$ and the image $I$, but also with a set of pre-computed, low-level hints $H$ (e.g., global brightness, sharpness, color metrics). The rater then outputs a single, binary veracity decision $D_{\text{HACF}}$ (i.e., “Pass” or “Fail”) for the entire sample: $D_{\text{HACF}}=\text{LLM}_{\text{Rater}}(R,I,H)$ We only retain samples with $D_{\text{HACF}}=\text{``Pass"}$, ensuring that only data with high-veracity reasoning is kept for training.
Figure 3: The GR-IQA dataset curation pipeline. It uses (1) Visual Reliance Filtering (VRF) to ensure visual grounding via token probabilities, and (2) Hint-Augmented Consistency Filtering (HACF) to perform sentence-level, hint-based checks for unfaithful text.
3.1.2 Grounded Rationale Fine-Tuning
Following the curation of our high-fidelity GR-IQA dataset, we proceed to the SFT stage. We fine-tune the VLM, parameterized by $\theta$, to auto-regressively generate the complete ground-truth response $S_{\text{structured}}=(R,A)$, where $R$ represents the textual rationale and $A$ denotes the answer. Given the image $I$ and the prompt $T_{\text{CoT}}$, the model is trained to predict the tokens of $S_{\text{structured}}$. This is achieved by minimizing the standard cross-entropy (CE) loss $\mathcal{L}_{\text{SFT}}$: $\mathcal{L}_{\text{SFT}}(\theta)=-\sum_{t=1}^{|S|}\log p_{\theta}(S_{t}\mid S_{<t},I,T_{\text{CoT}})$ where $S_{t}$ is the $t$-th token in the ground-truth sequence $S_{\text{structured}}$, and $|S|$ is the total length of the sequence.
3.2 Self-Guided Exploration
KL-Coverage Regularizer. Recent studies on applying Reinforcement Learning (RL) to large language models (LLMs) [13, 10] highlight a critical challenge: policy entropy often drops sharply at the onset of training, declining monotonically to near zero. This “entropy collapse” severely limits the model’s ability to explore, leading to performance plateaus. In the context of IQA, this issue is particularly detrimental. It leads to a collapse in the diversity of both reasoning paths and predicted rating scores. For instance, existing RL-based IQA methods, such as Visualquality-R1 [67], suffer from “score collapse.” On the KonIQ [18] test set, this method’s output unique score ratio is merely 2.04%, in stark contrast to the 71.34% of the ground-truth Mean Opinion Scores (MOS) distribution (when rounded to two decimal places). To address this problem, we propose the KL-Coverage regularizer. This approach is inspired by the use of KL penalties to constrain policy updates [47] and the recent finding that high covariance between action log-probabilities and logit changes leads to rapid policy entropy collapse [13]. Our regularizer is thus designed to specifically suppress numerical tokens that exhibit this high covariance.
Given a batch of $N$ rollout tokens, let $\pi_{\theta}(y_{i}|y_{<i})$ denote the policy’s probability for token $y_{i}$ given its prefix $y_{<i}$, and let $A(y_{i})$ be its associated advantage. We first compute the batch-level mean log-probability $\overline{\log\pi}$ and mean advantage $\overline{A}$: $\overline{\log\pi}=\frac{1}{N}\sum_{j=1}^{N}\log\pi_{\theta}(y_{j}|y_{<j}),\quad\overline{A}=\frac{1}{N}\sum_{j=1}^{N}A(y_{j}).$ We then define a token-wise covariance score $Cov(y_{i})$ as the centered cross-product:
$Cov(y_{i})=\left(\log\pi_{\theta}(y_{i}|y_{<i})-\overline{\log\pi}\right)\left(A(y_{i})-\overline{A}\right).$ (1)
Crucially, our regularizer mainly targets the numerical tokens responsible for the final score. We first define a candidate set $\mathcal{N}_{ans}$ comprising all numerical tokens within the <answer>...</answer> tags. We then rank the tokens in this candidate set $\mathcal{N}_{ans}$ by their $Cov(y_{i})$ scores. We define a binary mask $M_{i}$ for tokens $y_{i}\in\mathcal{N}_{ans}$, where $M_{i}=1$ if the token is in the top-$p$ proportion (e.g., $p=0.02$), and $M_{i}=0$ otherwise. Here, $p$ is a hyperparameter defining the fraction of these candidate tokens to be regularized.
Finally, we impose the KL penalty only on these selected tokens (where $M_{i}=1$). The KL-Coverage loss, $\mathcal{L}_{KLC}$, is computed as the mean KL divergence between the old policy $\pi_{\theta_{\text{old}}}$ and the current policy $\pi_{\theta}$, averaged only over these masked-in tokens:
$\mathcal{L}_{KLC}=\frac{\sum_{y_{i}\in\mathcal{N}_{ans}}M_{i}\cdot D_{KL}\left(\pi_{\theta_{\text{old}}}(y_{i}|y_{<i})\middle\|\pi_{\theta}(y_{i}|y_{<i})\right)}{\sum_{y_{i}\in\mathcal{N}_{ans}}M_{i}}.$ (2)
Progressive Re-sampling Strategy. Our training data suffers from a long-tailed score distribution, leading to poor performance on scarce score intervals (e.g., very high or low quality). To mitigate this data bias, we adopt a multi-stage re-sampling strategy. The model is first trained on the original data distribution, and in subsequent stages, we progressively increase the sampling frequency of these under-represented score intervals. This allows the model to first learn the general distribution and then fine-tune on rarer data, improving its robustness across the entire score range.
Figure 4: Qualitative comparison of Zoom-IQA with competing methods (Q-insight [29], VisualQuality-R1 [67]). We highlight: correct descriptions, incorrect descriptions, and the uncertainty-aware reasoning unique to our model.
Format Reward. This ensures that the model’s output strictly adheres to our required structured reasoning format. Specifically, the reasoning process, enclosed in <think>...</think> tags, must explicitly articulate key components such as "Directions for Improvement" and "Issues to Avoid". Furthermore, the final decision must be provided in a structured <answer>...</answer> tag, containing elements like bbox_2d and rating. If any of these formats are incorrect, the format score is 0. Only when all formats are correct can the model achieve the format score of 1.0 as defined:
[…]
Score Reward. This reward encourages the model to predict a quality rating $r_{\text{pred}}$ that is close to the ground-truth score $r_{\text{gt}}$. We define this as a continuous Gaussian reward based on their difference:
$R_{\text{score}}=\exp\left(-\frac{(r_{\text{pred}}-r_{\text{gt}})^{2}}{2\sigma^{2}}\right).$ (4)
where $\sigma$ is a hyperparameter controlling the sensitivity of the reward.
Rank Reward. To ensure the model learns relative quality ordering, we define a rank reward $R_{\text{rank}}(x_{i})$ based on pairwise comparisons within a batch, inspired by the Thurstone model [57]. It is computed as:
$R_{\text{rank}}(x_{i})=\frac{1}{B-1}\sum_{j\neq i}\left(\sqrt{\hat{p}_{ij}p^{*}_{ij}}+\sqrt{(1-\hat{p}_{ij})(1-p^{*}_{ij})}\right),$ (5)
where $p_{ij}=p(x_{i},x_{j})$ is the ground-truth preference derived from MOS, indicating if $\text{MOS}(x_{i})>\text{MOS}(x_{j})$. $p_{k,ij}=p_{k}(x_{i},x_{j})$ is the model’s predicted preference probability, calculated using the Thurstone model: $p_{k,ij}=\Phi\left(\frac{\mu_{i}-\mu_{j}}{\sqrt{v_{i}+v_{j}}}\right)$ Here, $\mu$ and $v$ represent the estimated mean and variance of the model’s rating distribution for an input, and $\Phi$ is the standard normal CDF.
Table 1: PLCC / SRCC comparison on the score regression tasks between our method and other competitive IQA methods. All methods except handcrafted ones are trained on the KonIQ dataset.
Category Methods KonIQ SPAQ KADID PIPAL LiveW AGIQA CSIQ
Handcrafted NIQE [40] 0.533 / 0.530 0.679 / 0.664 0.468 / 0.405 0.195 / 0.161 0.493 / 0.449 0.560 / 0.533 0.718 / 0.628
BRISQUE [39] 0.225 / 0.226 0.490 / 0.406 0.429 / 0.356 0.267 / 0.232 0.361 / 0.313 0.541 / 0.497 0.740 / 0.556
Non-VLM NIMA [56] 0.896 / 0.859 0.838 / 0.856 0.532 / 0.535 0.390 / 0.399 0.814 / 0.771 0.715 / 0.654 0.695 / 0.649
HyperIQA [53] 0.917 / 0.906 0.791 / 0.788 0.506 / 0.468 0.410 / 0.403 0.772 / 0.749 0.702 / 0.640 0.752 / 0.717
DBCNN [77] 0.884 / 0.875 0.812 / 0.806 0.497 / 0.484 0.384 / 0.381 0.773 / 0.755 0.730 / 0.641 0.586 / 0.572
MUSIQ [25] 0.924 / 0.929 0.868 / 0.863 0.575 / 0.556 0.431 / 0.431 0.789 / 0.830 0.722 / 0.630 0.771 / 0.710
ManIQA [70] 0.849 / 0.834 0.768 / 0.758 0.499 / 0.465 0.457 / 0.452 0.849 / 0.832 0.723 / 0.636 0.623 / 0.627
VLM (w/o & w/ reasoning) CLIP-IQA+ [59] 0.909 / 0.895 0.866 / 0.864 0.653 / 0.654 0.427 / 0.419 0.832 / 0.805 0.736 / 0.685 0.772 / 0.719
C2Score [81] 0.923 / 0.910 0.867 / 0.860 0.500 / 0.453 0.354 / 0.342 0.786 / 0.772 0.777 / 0.671 0.735 / 0.705
Q-Align [65] 0.941 / 0.940 0.886 / 0.887 0.674 / 0.684 0.403 / 0.419 0.853 / 0.860 0.772 / 0.735 0.671 / 0.737
DeQA [71] 0.953 / 0.941 0.895 / 0.896 0.694 / 0.687 0.472 / 0.478 0.892 / 0.879 0.809 / 0.729 0.787 / 0.744
Q-Insight [29] 0.918 / 0.895 0.903 / 0.903 0.702 / 0.702 0.458 / 0.435 0.870 / 0.839 0.816 / 0.766 0.685 / 0.640
VisualQuality-R1 [67] 0.910 / 0.896 0.889 / 0.892 0.703 / 0.712 0.451 / 0.441 0.856 / 0.827 0.817 / 0.760 0.768 / 0.707
Zoom-IQA (Ours) 0.938 / 0.922 0.902 / 0.900 0.701/ 0.700 0.468 / 0.465 0.887 / 0.870 0.816/ 0.765 0.797 / 0.754
Table 2: Quantitative results for image quality description. We evaluate using four metrics: Accuracy (Acc.), Reasonableness (Reason.), Completeness (Compl.), and Confidence (Conf.) using two closed-source VLM evaluators: Gemini-2.5-Flash and GPT-5-mini.
Method KonIQ SPAQ
Gemini-2.5-flash GPT-5-mini Gemini-2.5-flash GPT-5-mini
Acc. $\uparrow$ Reason. $\uparrow$ Compl. $\uparrow$ Conf. $\uparrow$ Acc. $\uparrow$ Reason. $\uparrow$ Compl. $\uparrow$ Conf. $\uparrow$ Acc. $\uparrow$ Reason. $\uparrow$ Compl. $\uparrow$ Conf. $\uparrow$ Acc. $\uparrow$ Reason. $\uparrow$ Compl. $\uparrow$ Conf. $\uparrow$
DepictQA [73] 5.40 5.49 5.51 7.96 4.54 5.09 4.41 7.80 6.04 6.39 6.14 7.86 4.80 5.08 4.46 6.87
VisualQuality-R1 [67] 7.29 7.60 7.29 7.57 6.10 6.05 6.02 6.79 8.32 8.35 7.70 7.55 6.61 6.67 5.51 6.82
Q-Insight [29] 7.17 7.44 7.08 6.93 5.32 5.74 5.29 6.15 7.84 8.02 7.51 6.98 6.22 6.35 5.54 5.90
Zoom-IQA(Ours) 8.72 8.80 8.30 8.60 6.93 6.93 6.61 7.98 8.63 8.69 8.47 8.63 6.97 7.27 6.79 7.99
4 Experiments
4.1 Experimental Settings
Implementation Details. We initialize Qwen2.5-VL-7b [2] as our base model during the first cold start stage, in which training is performed with a batch size of 2, 8 gradient accumulation steps, a learning rate of $2.5\times 10^{-6}$, and a warm-up ratio of 0.3. For GRPO, we train the finetuned model after the cold start stage with a batch size of 1, 2 gradient accumulation steps, a learning rate of $1\times 10^{-6}$, and a KL penalty coefficient of $\beta=0.04$. The number of generated responses N is set to 8.
Datasets and Metrics. For the first cold start stage, we applied SFT with our collected high-quality CoT datasets using cross-entropy loss. For the score regression task, we conduct training and evaluation on six IQA datasets grouped into three categories: (1) In-the-wild datasets, including KonIQ [18], SPAQ [15], and LIVE-Wild [16]; (2) Synthetic distortion datasets, including KADID [31], PIPAL [23], and CSIQ [26]; (3)AI-generated datasets, including AGIQA [28]. We adopt the pearson linear correlation coefficient (PLCC) and spearman rank-order correlation coefficient (SRCC) as metrics to evaluate performance on the score regression task, following previous works [71, 29].
4.2 Comparison and Evaluation
Image Quality Score Regression. We compare our method with SOTA IQA methods in three different categories: (I) handcrafted, including NIQE [40] and BRISQUE [39] ; (II) deep learning-based, NIMA [56], HyperIQA [53], DBCNN [77], MUSIQ [25], and ManIQA [70] (III) MLLM-based models, CLIP-IQA+ [60], C2Score [81], Q-Align [65], DeQA-Score [71], and Q-Insight [29]. Since VisualQualityR1 [67] did not report a KonIQ-only trained model, we retrained it with its official training code. As shown in Table 1, our approach achieves comparable performance compared with existing baselines across various synthetic and real-world benchmarks. When comparing with state-of-the-art IQA methods [29, 67] w/ reasoning capability, our Zoom-IQA presents consistently superior performance across almost all the benchmarks. Furthermore, a qualitative comparison demonstrates the superiority of our region-aware reasoning over competing methods (Fig. 4).
Table 3: Ablation studies on each component with PLCC / SRCC metrics. Models are trained on KonIQ.
SFT
Score
Reward
Rank
Reward
KL-Coverage
Regularizer Loss
Prog.
Training
KonIQ SPAQ KADID PIPAL LIVE-Wild AGIQA-3K CSIQ
1 0.836 / 0.806 0.849 / 0.839 0.632 / 0.621 0.431 / 0.426 0.806 / 0.762 0.766 / 0.697 0.688 / 0.645
2 0.906 / 0.887 0.892 / 0.884 0.709 / 0.699 0.450 / 0.447 0.854 / 0.827 0.784 / 0.739 0.772 / 0.710
3 0.928 / 0.915 0.896 / 0.892 0.677 / 0.664 0.379 / 0.392 0.875 / 0.865 0.807 / 0.745 0.715 / 0.696
4 0.908 / 0.890 0.888 / 0.882 0.683 / 0.669 0.455 / 0.446 0.874 / 0.847 0.806 / 0.738 0.759 / 0.718
5 0.932 / 0.918 0.898 / 0.895 0.665 / 0.652 0.458 / 0.455 0.881 / 0.865 0.810 / 0.758 0.791 / 0.749
6 0.938 / 0.922 0.902 / 0.900 0.701 / 0.700 0.468 / 0.465 0.887 / 0.870 0.816 / 0.765 0.797 / 0.754
Figure 5: Qualitative evaluation of reasoning quality on the image restoration task.
Image Quality Reasoning. To validate the effectiveness and accuracy of our reasoning chains, we follow common practices [7, 22] to employ a VLM-as-judge evaluation methodology on the KonIQ and SPAQ datasets. We utilize two powerful, closed-source VLMs (Gemini-2.5-Flash [12] and GPT-5-mini) as evaluators, which are tasked to score the generated descriptions on a 1-to-9 scale across four key criteria: Accuracy, Reasonableness, Completeness, and Confidence. To ground the assessments and ensure objectivity, the VLMs are prompted with the image, the generated reasoning chain, and corresponding low-level image indicators (e.g., brightness, sharpness) for cross-referencing. The detailed definitions of each metric and the full prompt structure are provided in the Appendix. As shown in Table 2, our method (Zoom-IQA) consistently and significantly outperforms all baselines [73, 67, 29] across both datasets and under the scrutiny of both VLM evaluators, indicating the superiority of our reasoning reliability.
Reasoning-guided Restoration. High-quality reasoning should be able to provide reliable guidance for downstream tasks such as restoration [62, 75, 61]. To further demonstrate the superiority of Zoom-IQA’s reasoning capability, we extract IQA reasoning as guidance for SUPIR [75], a state-of-the-art restoration model that accepts textual prompts. We replace its default prompt generator (LLaVA-1.5-13b [32]) and instead feed it the textual reasoning outputs from Q-Insight [29], VisualQuality-R1 [67], and our own Zoom-IQA to examine how IQA reasoning affects restoration outputs. As visualized in Figure 5, the reasoning generated by our method provides more appropriate and specific guidance, enabling SUPIR to restore fine-grained, detailed textures from the input images, which are overlooked or poorly reconstructed when guided by the reasoning from other methods. Such a comparison demonstrates the practical effectiveness of our model’s reasoning, highlighting the reliability of our IQA reasoning.
4.3 Ablation Study
We conduct a comprehensive ablation study to evaluate the contribution of each proposed component. The results, measured by PLCC and SRCC across seven commonly used benchmarks, are presented in Table 3. Our analysis begins with the SFT model (Row 1), which serves as the baseline.
Impact of Reward Signals. The comparison between Row 2 (Rank Reward only) and Row 3 (Score Reward only) indicates the effectiveness of Rank Reward for synthetic benchmarks (e.g., KADID: 0.709 vs. 0.677; CSIQ: 0.772 vs. 0.715). Besides, the Score Reward benefits real-world benchmarks (e.g., KonIQ: 0.928 vs. 0.906; SPAQ: 0.896 vs. 0.892). Such a comparison shows that both rewards contribute to the assessment of image quality.
Effect of KL-Coverage Regularizer Loss. We evaluate the impact of our KL-Coverage loss by comparing Row 4 (without KL) to Row 5 (with KL). Introducing such a regularizer brings consistent performance gains across the majority of datasets, such as on KonIQ (0.908 to 0.932), SPAQ (0.888 to 0.898), and CSIQ (0.759 to 0.791), indicating its effectiveness in preventing mode collapse and encouraging diverse reasoning.
Effect of Progressive Training. Comparing Row 5 (full model without) with Row 6 (full model with), progressive training provides a consistent performance lift across all datasets (e.g., KonIQ: 0.932 to 0.938; KADID: 0.665 to 0.701). Such an improvement stems from the strategy’s ability to effectively handle imbalanced quality data, i.e., conducting a superior evaluation on images at the extremes of the quality spectrum (i.e., very high or very low quality).
5 Conclusion
In this work, we proposed Zoom-IQA, a novel IQA framework that uses an iterative process of reasoning and zooming to focus on quality-relevant regions and generate accurate chain-of-thought reasoning. To encourage reliable reasoning, we first build the fine-grained Grounded-Rationale-IQA (GR-IQA) dataset. We further present key training strategies, including a two-stage scheme, reward designs, a KL-Coverage regularizer, and progressive resampling. We verify the effectiveness of our designs via extensive experiments from multiple aspects, including score prediction, reasoning examination, and the downstream application, together with a thorough ablation study. We believe our work could motivate future development in various domains, such as designing IQA data pipelines with automated data labeling, enhancing the reasoning reliability of IQA, and building more robust, interactive perceptual models.
Supplementary Material
6 Prompt Templates for GR-IQA Dataset Construction and Filtering
6.1 Data Construction Prompts
Tab. 4 and 5 detail the prompts used to construct our Grounded-Rationale-IQA (GR-IQA) dataset with Gemini-2.5-pro [12]. The thinking rationale comprises four components:
  1. 1.
    Image Quality Summary: A direct assessment of technical quality.
  2. 2.
    Directions for Improvement: An aspirational description of the ideal image.
  3. 3.
    Issues to Avoid: A detailed description of existing technical artifacts.
  4. 4.
    Decision & Rationale: A comprehensive justification covering the initial rating, crop analysis, and final decision.
A key feature of our design is that both the Directions for Improvement and Issues to Avoid are supported by regional findings. This spatial grounding allows the reasoning path to be effectively split into positive and negative prompts, boosting performance in downstream tasks.
Table 4: Prompt for Initial Image Quality Assessment (Stage 1). This prompt directs the model to act as an image quality expert, perform a preliminary evaluation, and decide whether a high-resolution crop is necessary to resolve uncertainties before making a final judgment.
Component Description
[…]
Persona The model is instructed to act as an Image Quality Expert.
[…]
Reasoning Sections
(in <think>)
The <think> block must contain exactly four labeled paragraphs: 1. Image Quality Summary: A concise verdict on technical flaws. 2. Directions for Improvement: Aspirational description of a perfect image (positive framing). 3. Issues to Avoid: Description of existing technical problems (negative framing). 4. Decision & Rationale: The initial rating, the crop/final decision, and its justification.
[…]
(in <answer>)
The <answer> block contains a JSON object with the following keys: "bbox_2d": Coordinates [x1, y1, x2, y2] for the crop, or [0,0,0,0] if no crop is needed. "rating": The initial quality score (e.g., 3.50). "tool": Either "crop" to request a zoom-in, or "final" to conclude the assessment.
Key Constraints - Reasoning must be in compact paragraphs, not bullet points. - Sections 2 and 3 must maintain strictly positive and negative language, respectively. - All feedback must be grounded in specific, named regions of the image.
Table 5: Prompt for Final Image Quality Assessment with Crop (Stage 2). This prompt is used after a crop has been generated in Stage 1. It instructs the model to synthesize information from the original image, the crop, and its own prior reasoning to produce a definitive and well-justified final quality score.
Component Description
[…]
Persona The model continues to act as an Image Quality Expert.
[…]
Reasoning Sections
(in <think>)
The <think> block is restructured to focus on the new evidence from the crop: 1. Crop Inspection Summary: A summary of what the crop confirmed or revealed. 2. Directions for Improvement: Aspirational goals based on details now visible within the crop. 3. Issues to Avoid: Technical problems confirmed or newly discovered within the crop. 4. Final Decision & Rationale: Explicitly references the initial rating and explains how the crop’s findings led to a rating upgrade, downgrade, or confirmation.
[…]
(in <answer>)
The JSON output is now always final: "bbox_2d": Always [0, 0, 0, 0]. "rating": The final, definitive quality score (e.g., 4.25). "tool": Always "final".
In-Context Learning The prompt includes three detailed examples demonstrating how to handle different scenarios: 1. Rating Downgrade: When the crop reveals the quality is worse than suspected. 2. Rating Upgrade: When the crop reveals the quality is better than suspected. 3. Rating Confirmation: When the crop confirms the initial assessment.
6.2 Visual Reliance Filtering
Due to the limited API access of closed-source VLMs, we cannot compute the log-probabilities for predefined text sequences (or specific candidate answers). Consequently, existing hallucination detection methods relying on offline contrastive probability analysis [68] are inapplicable. Furthermore, prior online decoding strategies [27, 20] typically inject image distortions to verify consistency. While effective for semantic-level tasks—where content identity remains robust to noise—this approach is fundamentally incompatible with IQA. Since IQA aims to precisely evaluate visual degradation, introducing artificial distortion alters the target attribute itself. This makes it difficult to disentangle whether the model is responding to the original image artifacts or the injected noise.
To address these limitations, we propose Visual Reliance Filtering (VRF). As illustrated in Fig. 6, VRF filters out samples where the model exhibits low visual dependency. We compare the VLM’s outputs under two distinct conditions: 1) conditioned on both the image $I$ and the textual rationale $R_{A}$ and 2) conditioned only on the partial rationale $R_{A}$ (without visual input). If the outputs are excessively similar, we discard the sample, as this indicates the visual input $I$ was non-essential and the response was driven primarily by language priors. In our experiments, we set the thresholds for rating difference, Bounding Box IoU, and entropy difference to 0.05, 0.5, and 0.01, respectively.
6.3 Hint-Augmented Consistency Filtering
Tab. 6 presents examples of data discarded by our Hint-Augmented Consistency Filtering with Qwen-2.5-32b [69]. Specifically, we leverage the low-level hints from KonIQ [18] and the quality hints from KonIQ++ [52] as reference; this allows our method to effectively filter out generated prompts that are inconsistent with human labels.
Table 6: Example of Filtered-Out Data in Hint-Augmented Consistency Filtering.
Example 1 2
Image [Uncaptioned image] [Uncaptioned image]
Generated prompt <think> 1) Image Quality Summary: The image suffers from significant digital noise, which is evident across the entire frame, especially in darker areas like the car and the person‘s jacket. The lighting is harsh due to direct on-camera flash, creating strong specular highlights on reflective surfaces and deep shadows, resulting in poor dynamic range. The overall focus is soft, lacking the crispness needed to resolve fine details. 2) Directions for Improvement: A higher quality image would render the “Cannondale” text on the bicycle frame with perfect sharpness and legibility. The person’s black jacket would appear with smooth, clean textures, and the silver fork of the bicycle would be detailed without harsh glare. The fine spokes of the wheels would be individually distinct and clear. 3) Issues to Avoid: There is excessive digital noise across the image, particularly degrading the texture on the car’s body and the person’s jacket. The direct flash causes blown-out specular highlights on the bicycle’s silver components and the car’s taillight. A general softness affects the entire scene, preventing sharp focus on key elements like the bicycle frame and the rider’s clothing. 4) Decision & Rationale: The image’s technical flaws, namely the pervasive noise, harsh lighting, and soft focus, are global and clearly visible without needing a closer look. The provided hint crops confirm these issues are present in key areas of detail. Therefore, a final rating can be given without cropping. The combination of these significant problems warrants a low rating, placing the image in the “Poor” category. </think> <answer> BBox 2D: [0, 0, 0, 0] Rating: 2.25 Tool: Final </answer> <think> 1) Image Quality Summary: The image quality is poor due to a combination of technical issues. It suffers from a soft focus across the entire frame, with no clear point of sharpness, and a very shallow depth of field. There is significant digital noise, particularly noticeable in the darker background areas, which gives the image a grainy texture. The lighting is harsh, likely from a direct flash, creating unflattering specular highlights on the flowers and vase and resulting in an overall flat appearance. 2) Directions for Improvement: A well-executed photograph would feature sharp focus on the central red roses, clearly defining the texture of their fabric petals. Diffused, soft lighting would elegantly highlight the brown and white patterns of the ribbed vase without creating harsh glare. Greater depth of field would render the delicate small white flowers with clarity, adding textural variety and depth to the floral arrangement. 3) Issues to Avoid: The image exhibits a pervasive softness that leaves the petals of the red roses and the details of the small white flowers blurry and undefined. Harsh, blown-out highlights from a direct flash are present on the shiny, iridescent ribbons within the bouquet and across the glossy surface of the vase. 4) Decision & Rationale: I will make a final decision as the image’s quality issues are immediately apparent and consistent throughout the frame. The softness, prominent noise, and harsh lighting are global problems that a crop would not help to clarify; it would only magnify these existing flaws. </think> <answer> BBox 2D: [0, 0, 0, 0] Rating: 1.75 Tool: Final </answer>
Low-level hint MOS $\uparrow$: 3.80 Brightness $\uparrow$: 0.2579 Contrast $\uparrow$: 0.1604 Colorfulness $\uparrow$: 0.3846 Sharpness $\uparrow$: 19.23 MOS $\uparrow$: 3.68 Brightness $\uparrow$: 0.2662 Contrast $\uparrow$: 0.1779 Colorfulness $\uparrow$: 0.0521 Sharpness $\uparrow$: 20.07
Quality hint QMOS $\uparrow$: 4.01 Artifacts $\downarrow$: 0.0000 Blurriness $\downarrow$: 0.033 QMOS $\uparrow$: 3.85 Artifacts $\downarrow$: 0.098 Blurriness $\downarrow$: 0.066
7 Image Quality Reasoning
Tab. 7 presents the prompts used to validate the effectiveness of our reasoning chains on the KonIQ and SPAQ [15] datasets. These datasets provide low-level image attributes (brightness, contrast, colorfulness, and sharpness for KonIQ; brightness, colorfulness, contrast, noisiness, and sharpness for SPAQ) alongside MOS scores, which ensures more precise evaluation judgments.
8 Reasoning-guided Restoration
To further demonstrate the efficacy of Zoom-IQA’s reasoning capability, we evaluate its performance on the task of text-guided image restoration. We employ the DreamClear [1] framework for this experiment, primarily because its T5 text encoder [45] supports a significantly longer context window than the CLIP encoder [44] used in SUPIR [75], thereby accommodating the detailed reasoning prompts generated by our method. Specifically, we adhere to the standard DreamClear pipeline, where initial restoration results from a lightweight network (SwinIR [30]) are input into VLMs to generate guidance. We conduct experiments under three distinct prompt settings. In addition to the standard captions from LLaVA-1.6-13b [32] used by the original DreamClear, we separately employ the reasoning content from VisualQuality-R1 and Q-Insight, and the “Directions for Improvement” derived from our method’s thinking process. As illustrated in Figs. 7-9, restoration guided by our method’s reasoning exhibits superior perceptual texture quality, particularly in complex regions such as facial features (Fig. 8) and fur (Fig. 9).
9 More Qualitative Comparison Results
We provide additional qualitative comparisons between our method (Zoom-IQA) and competing methods (Q-Insight [29], VisualQuality-R1 [67]) on real-world images, as illustrated in Fig. 10 through Fig. 13. These results demonstrate the superiority of our reasoning mechanism, which not only identifies specific distortions but also explicitly localizes the distorted objects/regions. Furthermore, in complex scenes (e.g., Fig. 10 and Fig. 11), our method employs interactive, region-aware reasoning: it first hypothesizes potential flaws (green text), then grounds them via adaptive cropping (orange text), and finally verifies the degradation (blue text). This hypothesize-and-verify loop ensures a comprehensive assessment. Conversely, in scenes with simpler compositions (e.g., Fig. 12 and Fig. 13), our method directly detects global distortions without performing cropping operations.
10 Experiment Settings
10.1 Total Rewards
The total reward $R_{\text{total}}$ for a trajectory comprises several components designed to guide the behavior of Zoom-IQA, formulated as:
$R_{\text{total}}=R_{\text{format}}+\alpha R_{\text{score}}+\beta R_{\text{rank}},$ (6)
where $\alpha$ and $\beta$ are coefficients that balance the relative importance of score prediction and rank consistency. In our experiments, we set $\alpha=1$ and $\beta=2$, while the parameter $\sigma$ in the score reward (Eq. 4 in main paper) is set to 0.35.
10.2 Reinforcement Learning Training of Self-Guided Exploration
In the Zoom-IQA framework, the VLM first processes the query and the original image. If the initial visual information is insufficient, the model outputs a ”crop” action directed at a specific region, which initiates a subsequent reasoning turn. A localized crop is then fed back into the VLM to generate a refined response. This design establishes an iterative interaction pipeline, enabling the model to progressively resolve ambiguities and refine its assessment. To optimize this multi-turn trajectory, we employ the GRPO [49] objective:
$\begin{split}&\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{y_{k}\}_{k=1}^{G}\sim\pi_{\text{old}}}\Bigg[\frac{1}{G}\sum_{k=1}^{G}\frac{1}{\sum_{t}V_{k,t}}\sum_{t=1}^{|y_{k}|}V_{k,t}\\ &\quad\cdot\min\Bigg(r_{k,t}(\theta)\hat{A}_{k,t},\operatorname{clip}\Bigg(r_{k,t}(\theta),1-\epsilon,1+\epsilon\Bigg)\hat{A}_{k,t}\Bigg)-\beta\mathcal{L}_{\text{KLC}}\Bigg],\end{split}$ (7)
where $y_{k}$ denotes the $k$-th generated response sequence in the group. The probability ratio is defined as $r_{k,t}(\theta)=\frac{\pi_{\theta}(y_{k,t}|q,y_{k,<t},\mathcal{I}_{<t})}{\pi_{\text{old}}(y_{k,t}|q,y_{k,<t},\mathcal{I}_{<t})}$, $\mathcal{L}_{\text{KLC}}$ is the KL penalty term from Eq. 2. The advantage $\hat{A}_{k,t}$ is computed by standardizing the total reward $R_{\text{total},k}$ using the mean and standard deviation of the group rewards $\{R_{\text{total},k}\}_{k=1}^{G}$. Crucially, we restrict the optimization scope to the model’s active generation, treating all user prompts and visual tool outputs as fixed context. Formally, we define a binary validity mask $V_{k,t}$ for each token in the trajectory: $V_{k,t}=1$ if $y_{k,t}$ is a token generated by the VLM (e.g., within the <think> or <answer> blocks), and $V_{k,t}=0$ if it corresponds to user queries or input tokens provided by the image-crop tool.
Table 7: Prompt for the “Image Quality Reasoning” Evaluation Strategy. This prompt configures a model to act as an expert human evaluator. It explicitly defines the priority of visual evidence over objective metrics and specifies the handling of strictly logical conclusions for both single and multi-round reasoning.
Component Description
[…]
Persona The model is instructed to act as an Expert Image Quality Evaluator, utilizing the provided image as the primary source of truth while treating objective metrics only as supporting technical references.
[…]
Input Context The evaluator is provided with: 1. [Model Response]: The text generated by the target model (assessing either direct reasoning or the final conclusion of a multi-round process). 2. [Reference Indicators]: Objective metrics for cross-referencing: MOS, Sharpness, Brightness, Contrast, and Colorfulness.
[…]
Figure 6: Example of filtered-out data in Visual Reliance Filtering.
Figure 7: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [29], VisualQuality-R1 [67]) and the baseline (LLaVA-1.6-13b [32]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
Figure 8: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [29], VisualQuality-R1 [67]) and the baseline (LLaVA-1.6-13b [32]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
Figure 9: Qualitative comparison of Zoom-IQA against competing methods (Q-Insight [29], VisualQuality-R1 [67]) and the baseline (LLaVA-1.6-13b [32]) on the image restoration task. (Upper) The text guidance generated by each method, with accurate descriptions highlighted. (Lower) The corresponding restored results utilizing these text prompts. Please zoom in for better details.
Figure 10: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [29] and VisualQuality-R1 [67]). We highlight correct descriptions and incorrect descriptions, in addition to the uncertainty-aware and verifying reasoning unique to our model. The bbox indicates a large, cropped zoom requested by Zoom-IQA, clearly showing the blurriness. While both Q-Insight and our method correctly identify both the blurriness and noise, VisualQuality-R1 only recognizes the blurriness and ignores the noise. Crucially, Q-Insight provides only general distortion information, whereas our method not only specifies the distortion types but also precisely identifies the object or region suffering from the distortion. Please zoom in for more details.
Figure 11: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [29] and VisualQuality-R1 [67]). We highlight correct descriptions and incorrect descriptions, along with the uncertainty-aware and verifying reasoning unique to our model. The bbox indicates a large, cropped zoom requested by Zoom-IQA, clearly showing the digital noise. While Zoom-IQA accurately identifies the issues, VisualQuality-R1 fails to detect minor focus problems and digital noise, claiming there is no significant noise and that details are distinguishable. Furthermore, although Q-Insight notices the softness problem, it reaches a contrastive conclusion (asserting the image has “reasonable detail” but “not a great amount of fine detail to point out”) and also fails to recognize the digital noise. Please zoom in for more details.
Figure 12: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [29], VisualQuality-R1 [67]). We highlight: correct descriptions and incorrect descriptions. The image exhibits minor digital artifacts, which was uniquely identified by Zoom-IQA. Please zoom in for more details.
Figure 13: Qualitative comparison of Zoom-IQA with competing methods (Q-Insight [29], VisualQuality-R1 [67]). We highlight: correct descriptions and incorrect descriptions. The image exhibits slight overexposure, which was uniquely identified by Zoom-IQA. Furthermore, VisualQuality-R1 incorrectly classified motion blur as out-of-focus blur. Please zoom in for more details.